# METAL: A Multi-Agent Framework for Chart Generation with Test-Time Scaling

### v2

@@ Abstract
Chart generation aims to generate code to produce charts satisfying the desired visual properties, e.g., texts, layout, color, and type. It has great potential to empower the automatic professional report generation in financial analysis, research presentation, education, and healthcare. In this work, we build a vision-language model (VLM) based multi-agent framework for effective automatic chart generation. Generating high-quality charts requires both strong visual design skills and precise coding capabilities that embed the desired visual properties into code. Such a complex multi-modal reasoning process is difficult for direct prompting of VLMs. To resolve these challenges, we propose METAL (Multi-agEnT frAmework with vision Language models for chart generation), a multi-agent framework that decomposes the task of chart generation into the iterative collaboration among specialized agents. METAL achieves a 5.2% improvement in F1 score over the current best result in the chart generation task. Additionally, METAL improves chart generation performance by 11.33% over Direct Prompting with LLaMA 3.2-11b. Furthermore, the METAL framework exhibits the phenomenon of test-time scaling: its performance increases monotonically as the logarithm of computational budget grows from $2^{9}$ to $2^{13}$ tokens.
@@ 1 Introduction
Figure 1: Direct prompting of current VLMs (e.g. GPT-4o) often fails to generate charts that accurately replicate reference charts, resulting in errors in structure, color, and text alignment. Our proposed approach, METAL, tackle this challenge with iterative refinement through generation, critique, and revision. Our experiments show that increasing the logarithm of test-time compute recurrence and token usage leads to improved accuracy.
@@
Furthermore, we have two key findings through in-depth analysis: (1) Test-time scaling in the METAL framework: Recent works suggest the potential of scaling the number of tokens (computational budget) to enhance the reasoning performance of LLMs (Snell et al., 2024; Muennighoff et al., 2025). Specifically, utilizing more tokens during inference may lead to improved performance. In this work, we found that there is a near-linear relationship between the performance and the logarithm of the computational budget in experiments. Specifically, the performance of METAL increases monotonically as the logarithm of the computational budget grows from $2^{9}$ to $2^{13}$ tokens. (2) Modality-tailored critiques enhance self-correction: We observe that explicitly separating different modalities during the critique process—such as visual evaluation and code analysis—substantially enhances the multimodal self-correction capabilities of VLMs. Ablation study shows that METAL with the separate-critique design achieves a 5.16% improvement over the single-critique baseline.
@@ 2.1 Chart Generation with VLMs
Figure 2: Overview of METAL: A multi-agents system that consists of four specialized agents working in an iterative pipeline: (1) Generation Agent creates initial Python code to reproduce the reference chart, (2) Visual Critique Agent identifies visual discrepancies between the generated and reference charts, (3) Code Critique Agent analyzes the code and provides specific improvement guidelines, and (4) Revision Agent modifies the code based on the critiques. The process iterates until either reaching the accuracy target or maximum attempts limit.
@@ 3.2 METAL
1:$y_{0}\leftarrow G(x_{\text{ref}})$triangleright Generate initial code
@@ 4.2 Experiment Results
Similarly, for the GPT base model, the baseline methods demonstrated high performance with average F1 scores ranging from 81.12% to 81.70%. However, METAL outperformed these methods by a considerable margin, achieving an average F1 score of 86.46%. Specifically, our method improved 5.2% in average over Direct Prompting.
@@ Results
As shown in Table 2, the full METAL achieves the highest average F1 score, outperforming all ablated variants. Specifically, when only the visual critique agent is used (METALV), the system obtains an average F1 score of 84.12%, while the variant using only the code critique agent (METALC) achieves 82.89%. The unified agent variant (METALS) yields the lowest average performance at 81.22%.
@@
METALV 83.43 82.57 77.57 93.69 84.12
METALC 82.35 80.90 76.69 91.93 82.89
METALS 80.26 78.88 74.50 89.82 81.22
METAL 86.31 84.17 79.86 95.50 86.38
@@ 5.3 Why METAL
Figure 5: Case study of METAL’s progressive refinement from initial generation to perfect accuracy. Starting from Round 0’s initial generation (60% color accuracy, 84% text accuracy), the system iteratively improves the output. In Round 1, the system identifies and corrects Y-axis scale issues and missing annotations, achieving 100% text accuracy. Round 2 refines the color representations of distributions, achieving perfect accuracy across all metrics.
@@ 6 Case Study
In Round 1, two specialized critique agents analyze the generated chart. The visual critique agent detects inconsistencies in axis scaling and missing annotations, while the code critique agent identifies the corresponding code-level issues (e.g. incorrect tick intervals and absent annotations ). Based on these critiques, the revision agent modifies the chart by adjusting the Y-axis scale and adding the missing annotation. These corrections result in a significant improvement, reaching perfect text accuracy, though color accuracy remains unchanged.
In Round 2, the critique agents further refine the chart. The visual critique agent highlights inaccuracies in the color assignments of distributions, noting that the generated chart does not precisely match the reference chart’s colors. The code critique agent pinpoints the exact color discrepancies in the code and provides specific RGB values for correction. The revision agent incorporates these insights, adjusting the color specifications in the code. This final revision achieves perfect alignment with the reference chart, with 100% accuracy across all evaluation metrics.

### v3

@@ Abstract
Chart generation aims to generate code to produce charts satisfying the desired visual properties, e.g., texts, layout, color, and type. It has great potential to empower the automatic professional report generation in financial analysis, research presentation, education, and healthcare. In this work, we build a vision-language model (VLM) based multi-agent framework for effective automatic chart generation. Generating high-quality charts requires both strong visual design skills and precise coding capabilities that embed the desired visual properties into code. Such a complex multi-modal reasoning process is difficult for direct prompting of VLMs. To resolve these challenges, we propose METAL (Multi-agEnT frAmework with vision Language models for chart generation), a multi-agent framework that decomposes the task of chart generation into the iterative collaboration among specialized agents. METAL achieves a 5.2% improvement in the F1 score over the current best result in the chart generation task. Additionally, METAL improves chart generation performance by 11.33% over Direct Prompting with LLaMA 3.2-11b. Furthermore, the METAL framework exhibits the phenomenon of test-time scaling: its performance increases monotonically as the logarithm of computational budget grows from $2^{9}$ to $2^{13}$ tokens.
@@ 1 Introduction
Figure 1: Direct prompting of current VLMs (e.g. GPT-4o) often fails to generate charts that accurately replicate reference charts, resulting in errors in structure, color, and text alignment. Our proposed approach, METAL, tackle this challenge with iterative refinement through generation, critique, and revision. Our experiments show that increasing the logarithm of test-time compute recurrence and token usage leads to improved performance.
@@
Furthermore, we have two key findings through in-depth analysis: (1) Test-time scaling in the METAL framework: Recent works suggest the potential of scaling the number of tokens (computational budget) to enhance the reasoning performance of LLMs (Snell et al., 2024; Muennighoff et al., 2025). Specifically, utilizing more tokens during inference may lead to improved performance. In this work, we found that there is a near-linear relationship between the performance and the logarithm of the computational budget in experiments. Specifically, the performance of METAL increases monotonically as the logarithm of the computational budget grows from $2^{9}$ to $2^{13}$ tokens. (2) Modality-tailored critiques enhance self-correction: We observe that explicitly separating different modalities during the critique process—such as visual evaluation and code analysis—substantially enhances the multimodal self-correction capabilities of VLMs. Ablation study shows that METAL with the separate-critique design achieves a 5.6% improvement over the single-critique baseline.
@@ 2.1 Chart Generation with VLMs
Figure 2: Overview of METAL: A multi-agents system that consists of four specialized agents working in an iterative pipeline: (1) Generation Agent creates initial Python code to reproduce the reference chart, (2) Visual Critique Agent identifies visual discrepancies between the generated and reference charts, (3) Code Critique Agent analyzes the code and provides specific improvement guidelines, and (4) Revision Agent modifies the code based on the critiques. The process iterates until either reaching the verification score or maximum attempts limit.
@@ 3.2 METAL
1:$y_{0}\leftarrow G(x_{\text{ref}})$
@@ 4.2 Experiment Results
Similarly, for the GPT base model, the baseline methods demonstrated high performance with average F1 scores ranging from 81.12% to 82.32%. However, METAL outperformed these methods by a considerable margin, achieving an average F1 score of 86.46%. Specifically, our method improved 5.2% in average over Direct Prompting.
@@ Results
As shown in Table 2, the full METAL achieves the highest average F1 score, outperforming all ablated variants. Specifically, when only the visual critique agent is used (METALV), the system obtains an average F1 score of 84.31%, while the variant using only the code critique agent (METALC) achieves 82.96%. The unified agent variant (METALS) yields the lowest average performance.
@@
METALV 83.43 82.57 77.57 93.69 84.31
METALC 82.35 80.90 76.69 91.93 82.96
METALS 80.26 78.88 74.50 89.82 80.86
METAL 86.31 84.17 79.86 95.50 86.46
@@ 5.3 Why METAL
Figure 5: Case study of METAL’s progressive refinement from initial generation to perfect. Starting from Round 0’s initial generation (60% color score , 84% text score), the system iteratively improves the output. In Round 1, the system identifies and corrects Y-axis scale issues and missing annotations, achieving 100% text score. Round 2 refines the color representations of distributions, achieving perfect F1 score across all metrics.
@@ 6 Case Study
In Round 1, two specialized critique agents analyze the generated chart. The visual critique agent detects inconsistencies in axis scaling and missing annotations, while the code critique agent identifies the corresponding code-level issues (e.g. incorrect tick intervals and absent annotations ). Based on these critiques, the revision agent modifies the chart by adjusting the Y-axis scale and adding the missing annotation. These corrections result in a significant improvement, reaching perfect text score, though color score remains unchanged.
In Round 2, the critique agents further refine the chart. The visual critique agent highlights inaccuracies in the color assignments of distributions, noting that the generated chart does not precisely match the reference chart’s colors. The code critique agent pinpoints the exact color discrepancies in the code and provides specific RGB values for correction. The revision agent incorporates these insights, adjusting the color specifications in the code. This final revision achieves perfect alignment with the reference chart, with 100% f1 score across all evaluation metrics.